# Catch Causal Signals from Edges for Label Imbalance in Graph Classification

### v1

@@ Abstract
Despite significant advancements in causal research on graphs and its application to cracking label imbalance, the role of edge features in detecting the causal effects within graphs has been largely overlooked, leaving existing methods with untapped potential for further performance gains. In this paper, we enhance the causal attention mechanism through effectively leveraging edge information to disentangle the causal subgraph from the original graph, as well as further utilizing edge features to reshape graph representations. Capturing more comprehensive causal signals, our design leads to improved performance on graph classification tasks with label imbalance issues. We evaluate our approach on real-word datasets PTC, Tox21, and ogbg-molhiv, observing improvements over baselines. Overall, we highlight the importance of edge features in graph causal detection and provide a promising direction for addressing label imbalance challenges in graph-level tasks. The model implementation details and the codes are available on https://github.com/fengrui-z/ECAL.
@@ II-A Task formulation
For a general graph $G=(\mathcal{V},\mathcal{E})$ composed of the node set $\mathcal{V}$ and the edge collection $\mathcal{E}$, we represent it as a triplet of matrices $(A,X,E)$, where:
@@
  • $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ is the edge feature matrix, whose rows $E_{ij}$’s respectively denote the $d_{e}$-dimensional feature vector for the edge connecting nodes $i$ and $j$ in the graph.
@@ III-A Incorporating edge features into attention
Regarding the implementation of integrating edge features into the attention mechanism, it is intuitive to use the Edge-Featured Graph Attention Network (EGAT) [27]. In this module, the causality attention score $\alpha_{i}$ for each node $i$ and causality attention score $\alpha_{ij}$ for each edge $(i,j)$ is computed by considering both node and edge information. (We note this module, referred to as EGATv1, will shortly be enhanced and ultimately not employed in step ❶; however, this module will be evaluated in step ❷ graph representation.)
The node features $X_{i},X_{j}$ and the edge feature $E_{ij}$ are first transformed through linear projections; the concatenation of the transformed features can be expressed as:
$\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel E_{ij}^{\prime}\right],$
where $X_{i}^{\prime}=W_{x}^{(i)}X_{i}$, $X_{j}^{\prime}=W_{x}^{(j)}X_{j}$, and $E_{ij}^{\prime}=W_{e}E_{ij}$. The node embedding from the EGAT is computed as
@@
where the attention score $a_{ij}$ involves the comprehensive information from $\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel E_{ij}^{\prime}\right]$, computed as
$a_{ij}=\frac{\exp\left(\text{\scriptsize LeakyReLU}\left(a^{T}\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel E_{ij}^{\prime}\right]\right)\right)}{\sum_{k\in\mathcal{N}(i)\cup\{i\}}\exp\left(\text{\scriptsize LeakyReLU}\left(a^{T}\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel E_{ij}^{\prime}\right]\right)\right)},$
where $\mathcal{N}(i)$ is the neighborhood of node $i$ and $a$ indicates learnable parameters. Then we adopt two MLPs to estimate the node causality attention scores $\alpha_{i}$ and the edge causality attention scores $\alpha_{ij}$ respectively based on $X_{i}^{out}$ and $E_{ij}^{\prime}$:
@@
$\alpha_{ij},1-\alpha_{ij}$ $=\sigma(\mathrm{MLP}_{\mathrm{edge}}(E_{ij}^{\prime}))$ (3)
@@
An enhanced EGAT alternative: we note the edge causality attention scores in Equation 3 are computed using edge information $E_{ij}^{\prime}$, which is isolated from the neighborhood information. To better leverage edge features, we use another EGAT design [30], and we call it EGATv2 to distinguish it from the design above. This structure updates both edge and node features, allowing for a more effective capture of the dynamic relationships between nodes; in EGATv2, given the feature $\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel E_{ij}^{\prime}\right]$, the edge embedding is first calculated as ($W_{e}^{out}$ indicates learnable parameters):
$E_{ij}^{out}=\text{\scriptsize LeakyReLU}\left(W_{e}^{out}\left[X_{i}^{\prime}\parallel E_{ij}^{\prime}\parallel X_{j}^{\prime}\right]\right).$
The node embedding is next generated based on $E_{ij}^{out}$:
$X_{i}^{out}=\text{MLP}\left(\sum_{j\in\mathcal{N}(i)}a_{ij}\left(E_{ij}^{out}\right)X_{j}^{\prime}\right),$ (4)
where $a_{ij}$ is now a function of $E_{ij}^{out}$ and reads
$a_{ij}=\frac{\exp\left(a^{T}E_{ij}^{out}\right)}{\sum_{k\in\mathcal{N}(i)\cup\{i\}}\exp\left(a^{T}E_{ij}^{out}\right)}.$
We then follow a process akin to EGATv1, utilizing two MLPs to estimate the node causality attention scores and edge causality attention scores, with $X_{i}^{out},E_{ij}^{out}$ as inputs this time:
@@
$\alpha_{ij},1-\alpha_{ij}$ $=\sigma(\mathrm{MLP}_{\mathrm{edge}}(E_{ij}^{out}))$
where $\sigma(\cdot)$ is the softmax function. Compared to EGATv1, the edge causality attention scores are now estimated by the more comprehensive edge embedding $E_{ij}^{out}$. We term this design (with EGATv2) for step ❶ attention estimation as edge-enhanced causal attention learning (ECAL) along this paper.
@@ III-B1 Causal subgraph splitting
$G_{c}=(A,M_{x}\cdot X,M_{e}\cdot E),$ (5)
$G_{t}=(A,\bar{M_{x}}\cdot X,\bar{M_{e}}\cdot E),$ (6)
@@ III-B2 Edge-featured causal and trivial representations
$G_{c}=\{A,X^{(c)},E^{(c)}\},\quad G_{t}=\{A,X^{(t)},E^{(t)}\}.$ (7)
@@
$\hat{y}^{(c)}$ $=\Phi_{c}(h^{(c)}),\text{with~{}}h^{(c)}=f_{\text{readout}}^{(c)}(\text{EGAT}_{c}(X^{(c)},E^{(c)})),$
$\hat{y}^{(t)}$ $=\Phi_{t}(h^{(t)}),\text{with~{}}h^{(t)}=f_{\text{readout}}^{(t)}(\text{EGAT}_{t}(X^{(t)},E^{(t)})).$
@@ V Conclusions
This paper proposes an edge-enhanced framework for causal attention mechanisms, leveraging edge-featured graph attention networks to better separate causal and non-causal subgraphs while extracting richer graph representations. This design extends the capabilities of causal attention mechanisms, providing a flexible approach for diverse graph-level tasks in real-world applications.We validate our models on benchmark datasets (tailored to exhibit label imbalance), including PTC, Tox21, and ogbg-molhiv; our proposed methods outperform existing ones in label imbalance scenarios and showcase the necessity of incorporating edge features.

### v2

@@ Abstract
Despite significant advancements in causal research on graphs and its application to cracking label imbalance, the role of edge features in detecting the causal effects within graphs has been largely overlooked, leaving existing methods with untapped potential for further performance gains. In this paper, we enhance the causal attention mechanism through effectively leveraging edge information to disentangle the causal subgraph from the original graph, as well as further utilizing edge features to reshape graph representations. Capturing more comprehensive causal signals, our design leads to improved performance on graph classification tasks with label imbalance issues. We evaluate our approach on real-word datasets PTC, Tox21, and ogbg-molhiv, observing improvements over baselines. Overall, we highlight the importance of edge features in graph causal detection and provide a promising direction for addressing label imbalance challenges in graph-level tasks. The preprint version is available at https://arxiv.org/abs/2501.01707. The model implementation details and the codes are available on https://github.com/fengrui-z/ECAL.
@@ II-A Task formulation
For a general graph $G=(\mathcal{V},\mathcal{E})$ composed of the node set $\mathcal{V}$ and the edge collection $\mathcal{E}$, we represent it as a triplet of matrices $(\mathbf{A},\mathbf{X},\mathbf{E})$, where:
@@
  • $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ is the edge feature matrix, where for $(i,j)\in\mathcal{E}$, $\mathbf{e}_{ij}\in\mathbb{R}^{d_{e}}$ denotes the $d_{e}$-dimensional feature vector for the edge connecting nodes $i$ and $j$ in the graph.
@@ III-A Incorporating edge features into attention
Regarding the integration of edge features into the attention mechanism, it is intuitive to use the Edge-Featured Graph Attention Network (EGAT) [27]. In this module, the causality attention score $\alpha_{i}$ for each node $i$ and the causality attention score $\alpha_{ij}$ for each edge $(i,j)\in\mathcal{E}$ (i.e., node $j$ must be within $\mathcal{N}(i)$, the neighborhood of node $i$) are calculated considering the information from the node and the edge. (We note this module, referred to as EGATv1, will shortly be enhanced and ultimately not employed in step ❶; however, this module will be evaluated in step ❷ graph representation.)
The node features $X_{i},X_{j}$ and the edge feature $\mathbf{e}_{ij}$ are first transformed through linear projections; the concatenation of the transformed features can be expressed as:
$\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel\mathbf{e}_{ij}^{\prime}\right],$
where $X_{i}^{\prime}=W_{x}^{(i)}X_{i}$, $X_{j}^{\prime}=W_{x}^{(j)}X_{j}$, and $\mathbf{e}_{ij}^{\prime}=W_{e}\mathbf{e}_{ij}$. The node embedding from the EGAT is computed as
@@
where the attention score $a_{ij}$ involves the comprehensive information from $\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel\mathbf{e}_{ij}^{\prime}\right]$, computed as
$a_{ij}=\frac{\exp\left(\text{\scriptsize LeakyReLU}\left(a^{T}\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel\mathbf{e}_{ij}^{\prime}\right]\right)\right)}{\sum_{k\in\mathcal{N}(i)}\exp\left(\text{\scriptsize LeakyReLU}\left(a^{T}\left[X_{i}^{\prime}\parallel X_{k}^{\prime}\parallel\mathbf{e}_{ik}^{\prime}\right]\right)\right)},$
where $a$ indicates learnable parameters. Then we adopt two MLPs to estimate the node causality attention scores $\alpha_{i}$ and the edge causality attention scores $\alpha_{ij}$ respectively based on $X_{i}^{out}$ and $\mathbf{e}_{ij}^{\prime}$:
@@
$\alpha_{ij},1-\alpha_{ij}$ $=\sigma(\mathrm{MLP}_{\mathrm{edge}}(\mathbf{e}_{ij}^{\prime}))$ (3)
@@
An enhanced EGAT alternative: we note the edge causality attention scores in Equation 3 are computed using edge information $\mathbf{e}_{ij}^{\prime}$, which is isolated from the neighborhood information. To better leverage edge features, we use another EGAT design [30], and we call it EGATv2 to distinguish it from the design above. This structure updates both edge and node features, allowing for a more effective capture of the dynamic relationships between nodes; in EGATv2, given the feature $\left[X_{i}^{\prime}\parallel X_{j}^{\prime}\parallel\mathbf{e}_{ij}^{\prime}\right]$, the edge embedding is first calculated as ($W_{e}^{out}$ indicates learnable parameters):
$\mathbf{e}_{ij}^{out}=\text{\scriptsize LeakyReLU}\left(W_{e}^{out}\left[X_{i}^{\prime}\parallel\mathbf{e}_{ij}^{\prime}\parallel X_{j}^{\prime}\right]\right).$
The node embedding is next generated based on $\mathbf{e}_{ij}^{out}$:
$X_{i}^{out}=\text{MLP}\left(\sum_{j\in\mathcal{N}(i)}a_{ij}\left(\mathbf{e}_{ij}^{out}\right)X_{j}^{\prime}\right),$ (4)
where $a_{ij}$ is now a function of $\mathbf{e}_{ij}^{out}$ and reads
$a_{ij}=\frac{\exp\left(a^{T}\mathbf{e}_{ij}^{out}\right)}{\sum_{k\in\mathcal{N}(i)}\exp\left(a^{T}\mathbf{e}_{ik}^{out}\right)}.$
We then follow a process akin to EGATv1, utilizing two MLPs to estimate the node causality attention scores and edge causality attention scores, with $X_{i}^{out},\mathbf{e}_{ij}^{out}$ as inputs this time:
@@
$\alpha_{ij},1-\alpha_{ij}$ $=\sigma(\mathrm{MLP}_{\mathrm{edge}}(\mathbf{e}_{ij}^{out}))$
where $\sigma(\cdot)$ is the softmax function. Compared to EGATv1, the edge causality attention scores are now estimated by the more comprehensive edge embedding $\mathbf{e}_{ij}^{out}$. We term this design (with EGATv2) for step ❶ attention estimation as edge-enhanced causal attention learning (ECAL) along this paper.
@@ III-B1 Causal subgraph splitting
$G_{c}=(\mathbf{A},M_{x}\cdot\mathbf{X},M_{e}\cdot\mathbf{E}),$ (5)
$G_{t}=(\mathbf{A},\bar{M_{x}}\cdot\mathbf{X},\bar{M_{e}}\cdot\mathbf{E}),$ (6)
@@ III-B2 Edge-featured causal and trivial representations
$G_{c}=\{\mathbf{A},\mathbf{X}^{(c)},\mathbf{E}^{(c)}\},\quad G_{t}=\{\mathbf{A},\mathbf{X}^{(t)},\mathbf{E}^{(t)}\}.$ (7)
@@
$\hat{y}^{(c)}$ $=\Phi_{c}(h^{(c)}),\text{with~{}}h^{(c)}=f_{\text{readout}}^{(c)}(\text{EGAT}_{c}(\mathbf{X}^{(c)},\mathbf{E}^{(c)})),$
$\hat{y}^{(t)}$ $=\Phi_{t}(h^{(t)}),\text{with~{}}h^{(t)}=f_{\text{readout}}^{(t)}(\text{EGAT}_{t}(\mathbf{X}^{(t)},\mathbf{E}^{(t)})).$
@@ V Conclusions
In this paper, we proposed an edge-enhanced framework for causal attention mechanism, which applies edge-featured graph attention networks to better separate causal and non-causal subgraphs and extract the graph representation. This new design extends the capabilities of causal attention mechanisms, offering a promising direction for tackling diverse graph-level tasks in real-world applications. We validate the effectiveness of our proposed models on benchmark datasets (label imbalanced), including PTC, Tox21, and ogbg-molhiv; our proposed methods demonstrate superiority in label imbalance scenarios compared to existing methods and showcase the necessity of incorporating edge features.
Acknowledgments
This work is supported by SF Tech. Yifan Chen is supported by the Research Grants Council (RGC) under grant ECS-22303424, and GuangDong Basic and Applied Basic Research Foundation through the general project grant.